# Mitigating the Influence of Distractor Tasks in LMs with Prior-Aware Decoding

### v2

@@ 1 Introduction
Language models (LMs) have come to occupy a central role in a wide range of tasks, from data processing to the creation of instruction-following assistants. These models seem to both increase in performance and also develop new capabilities as they scale up in parameters, especially when given examples of a task (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). They see both widespread public use, and increasing integration into sensitive tasks where their versatile capabilities are key (Javaid et al., 2022). But despite their capabilities, language models have some unique limitations.
One such limitation are distractor tasks, where the models infer unintended secondary tasks from the prompt in addition to the intended one, leading to unwanted outputs(Wei et al., 2023). Distractor tasks range from simple ones such as completing a common sequence of tokens despite earlier instruction, prompt injections overriding the original instructions, to more complex and abstract ones, such as ensuring adherence to any previously observed directives. For example, the setting of a prompt injection attack can be characterised as two competing tasks: 1) the distractor task – following the most recent injected instructions, and 2) the intended task – following the initial instruction to ignore any injected instructions (Wei et al., 2023; Greshake et al., 2023; McKenzie et al., 2023).
One proposed cause of distractor tasks is that language models exhibit strong priors (McKenzie et al., 2023), where models learn certain patterns from the data so strongly that when the model is presented with a prompt that instigates the pattern, the tendency to continue the pattern distracts the model from explicit instructions. Even cutting-edge models struggle with seemingly trivial tasks in the presence of very common patterns. For example, as of May 2024 GPT-4o, when prompted with "Write all the numbers up to 30 except multiples of 13, and nothing else", will correctly exclude 13 but will fail to exclude 26.
@@
Other examples of such distractor tasks include repeating common misconceptions instead of true answers (Lin et al., 2021), and struggling to answer mathematical questions when the correct answer is very similar to an incorrect but a much more common pattern (Collins et al., 2023).
@@
Furthermore, for some tasks of this type, larger models actually perform worse: similar models with more parameters can be more likely to repeat common misconceptions and succumb to prompt injection attacks (McKenzie et al., 2023; Lin et al., 2021). This phenomenon, known as ‘inverse scaling’ (Lin et al., 2021), reveals a pressing need to better understand how these problems arise and how they can be fixed. Unlike other problems with LMs which are usually resolved by scaling up the model and training data size, the existing problems might become even more pronounced, and other similar issues might emerge.
The more integrated LMs become into modern society, the more crucial it is that we understand these phenomena and how they can be mitigated. It is important to be able to ensure that as models become more powerful they do not also become easier to subvert from their original instructions and into performing harmful tasks or disseminating dangerous information.
@@
Figure 1: Illustration of our geometric mixture model, also referred to as product of experts, and Prior-Aware Decoding on an instance of local strong prior in a task requiring a modification of a very common sequence.
One recent approach to deal with limitations of LMs and guide them towards desired outputs is contrastive decoding (Li et al., 2023b). Their method works at inference time and for a given prompt obtains two output token distributions: one using the original model, and one using a similar smaller model. It then contrasts the logits from the smaller model in the direction of the original logits, emphasizing the completions that the larger model prefers more than the small model, approximating a hypothetical model that is larger than the original model. However, in the context of mitigating the influence of distractor tasks this method might not be suitable given that these tasks can exhibit inverse scaling, which would make larger models even more susceptible.
Other similar contrastive inference methods methods have been proposed that instead of contrasting against a smaller model, use other undesirable logit distributions (Kim et al., 2023; Shi et al., 2023), or use extrapolation rather than simple contrasting. Our work can be similarly characterised as reducing the undesirable influence of distractor tasks.
Language models as product of experts. In order to describe the problem of distractor tasks more precisely, we characterise language models as product of experts Hinton (1999) which combine the predictions of several component models, each component model predicting the next token assuming a particular task or context. The component models here are proposed as an abstraction over possible model skills and behaviors – for instance one component model may perform the intended task while another component may perform the distractor task. It is then their relative weight in the joint model that determines which prediction prevails. See Figure 1 for an illustration of our model on a simple task of modifying a very common sequence.
Note that while we do not assume the components to be in any particular way represented within the model’s weights or activations, the Linear Representation Hypothesis (Jiang et al., 2024) hints at the possibility of such a representation, and offers a direction for future research. Another indirect evidence in this direction is the observed specialization of attention heads and their selection based on the context and given task (Li et al., 2023a).
We use the product of experts model view of LLMs as the primary model of the phenomenon of distractor tasks, modelling failures related to distractor tasks as an unintended component model dominating in the mixture of components and behaviors. Specifically, we frame it as a problem of strong priors in the domain of these components, where a component corresponding to a distractor task has a large prior probability.
Following this underlying model, we propose a general method of re-weighing component models via extrapolation between different geometric mixtures. In this paper we propose and demonstrate a concrete method of eliciting an undesirable mixture and extrapolating away from it.
Prior-Aware Decoding. Based on this framework, we introduce Prior-Aware Decoding (PAD) – a method to mitigate the influence of distractor tasks. We deliberately create a weakened prompt that is designed to make the model more susceptible to the distractor task. The prompt is weakened by removing information relating to the intended task, while retaining the parts responsible for the distractor task.
We then extrapolate the logits $L_{O}$ from the original prompt with logits $L_{W}$ from the weakened prompt via a linear combination $L_{\alpha}=L_{O}+\alpha\cdot(L_{O}-L_{W})$, with $\alpha$ being a tunable parameter. We then sample the tokens according to $L_{\alpha}$ using the standard Boltzmann (or softmax) distribution.
The technique requires two choices: selection of the weakened prompt, and the coefficient $\alpha$. We offer consistent approaches to generating the weakened prompts for each of our datasets and also propose a general method for prompt weakened prompt construction. We explore a range of values for $\alpha$. Crucially, this method does not require retraining: it operates at inference time by making two queries instead of one and computing a linear combination of the resulting logits.
@@
Our primary baseline is the unmodified completion according to the same model. Across the 44 model-task combinations we find that our technique outperforms the baseline on 41 combinations, the median absolute increase in proportion of tasks completed being 40%. See Figure 2 and Section 6 for the results.
The improvements of accuracy due to our technique suggests that viewing LMs as product of experts models can be a useful framework for modelling the behaviour of these models. Our research provides inroads for utilising capabilities which are dormant in models and which have been hard to elicit with prompting alone. This includes capabilities of immediate practical importance, such as resistance to prompt injection attacks, as well as understanding how to correct for biases learned in the training data.
@@
  • Interpreting language models as geometric mixtures (i.e. product of experts), we develop a model of a task selection mechanism to explain how distractor tasks cause poor performance on certain tasks.
@@ 2 Related Work
Inverse Scaling. For most tasks performance scales together with the model scale (model and dataset size, amount of compute). However, for certain tasks a phenomenon of inverse scaling is observed, where increasing the scale of the model leads to worse performance. It was introduced by Lin et al. (2021) to characterise the tendency of LMs to repeat common misconceptions. McKenzie et al. (2023) further developed the notion and provided the Inverse Scaling dataset as well as several informal categories. Wei et al. (2023) adds to the discussion, showing that of the original eleven tasks, only four persist in demonstrating inverse scaling on models of up to 540b parameters. We apply our method to three of those four tasks. Our work is the first to offer a formal account of the mechanisms that have so far only been informally hypothesised, and we use this to develop a novel technique that improves performance.
@@
Problems Related to Strong Priors. While in this paper we investigate the problem of strong priors in the domain of language modelling, a somewhat similar problem also occurs in diffusion models for image generation. Diffusion models have been observed to sometimes struggle to generate images that have a precise correspondence to the prompt. This problem is especially pronounced for complex prompts that include rare words, or describe implausible scenes, spatial relations, or compositionality. One technique for mitigating this problem in diffusion models is classifier-free guidance (Ho & Salimans, 2022). It samples the diffusion model using a linear combination of conditioned and unconditioned score estimates $(1+w)\cdot e(z_{t},c)-w\cdot e(z_{t})$, where $z_{t}$ is the image at the denoising step $t$, $c$ is the conditioning information such as the prompt, $e$ is the model and $w$ is the guidance weight. High guidance weight improves text-image alignment (Saharia et al., 2022), although high values often lead to over-saturated images or have a generally unnatural appearance. The technique that we present in this paper uses a somewhat similar combination of differently conditioned model outputs to mitigate the problem of strong priors in LMs.
Steering Language Models. Improving instruction-following is a special case of the more general research direction of controlling the behaviour of LMs. There are various techniques that have been proposed for it, such as fine-tuning on datasets exhibiting the desired behaviour (Dai & Le, 2015) or reinforcement learning from human feedback (Ziegler et al., 2019). Recently, Turner et al. (2023) proposed ‘activation addition’ for steering the behaviour of LMs without optimisation. For example, they bias the model towards affectionate outputs for a given prompt by adding activations at a certain layer from the same model when it was run on the prompt ‘love’. Like our approach, it works at inference time by combining the results of a model run on specifically selected prompts.
@@
Surface Form Competition and DC-PMI. Holtzman et al. (2021) coin ‘surface form competition’ as a phenomenon where different tokens representing the same concept compete for probability. For instance, the terms “computer” and “PC” represent the same concept but are different surface forms. Since the probability mass is finite, this competition lowers the probability of the correct answer. Domain Conditional Pointwise Mutual Information (DC-PMI) addresses this issue by reweighing each option according to its likelihood within the context of a specific task. This approach yields gains in zero-shot performance for a variety of GPT-2 and GPT-3 models on a selection of multiple-choice datasets. Our work applies a similar kind of technique - reweighing based on likelihood - but we do so to address a different set of tasks, including those which exhibit inverse scaling.
@@
Contrastive Inference Methods. One recent approach to deal with limitations of LMs and guide them towards desired outputs is contrastive decoding, introduced for language models by (Li et al., 2023b). Their method works at inference time and for a given prompt obtains two output token distributions: one using the original model, and one using a similar smaller model. It then contrasts the logits from the two models, selecting the token with the largest difference between the original ($L_{O}$) and smaller model ($L_{S}$), i.e. $\max_{t\in T^{\prime}}(L_{O}(t)-L_{S}(t))$ (limited to a subset $T^{\prime}$ of top tokens to mitigate noise in low-likelihood tokens).
However, in the context of mitigating the influence of distractor tasks, this method might not be suitable given that these tasks can exhibit inverse scaling, which would make larger models even more susceptible, as we demonstrate in Section 6.
Other similar contrastive inference methods have been proposed that instead of contrasting against a smaller model use other undesirable logit distributions ($L_{U}$) (Kim et al., 2023; Shi et al., 2023) based on providing context or contrasting to noisy instructions. (Kim et al., 2023) use another way of computing the contrast between the distributions: $\max_{t\in T^{\prime}}(L_{O}(t)-\varepsilon L_{U}(t))$ (with a limited set of tokens $T^{\prime}$). The linear extrapolation $\max_{t\in T}((1+\alpha)L_{O}(t)-\alpha L_{S}(t))$ of (Shi et al., 2023) is functionally equivalent to our method, applied in a different context.
In this work, we describe a formal model (Section 4) that provides a principled foundation for the contrastive inference method in the context of distractor tasks. The method reduces the undesirable influence of distractor tasks by extrapolating between the original $(L_{O})$ and weakened model logits $(L_{W})$ as $\max_{t\in T}(L_{O}(t)+\alpha(L_{O}(T)-L_{W}(t)))$.
While our logit extrapolation formula is functionally equivalent to the one used in (Shi et al., 2023), our method specifically targets distractor tasks, inverse scaling, and strong priors. By focusing on these aspects and providing a formal model of the scenario, our method offers a targeted and theoretically grounded approach to mitigating the influence of strong priors in language models.
Further work may be required for a joint generalization of the above approaches and a common treatment of the considered contexts. In contrast to the contrastive decoding of (Li et al., 2023b), contrastive decoding via logit extrapolation also has the advantage of not requiring a cutoff for the tokens considered, since even tokens with a very large contrast $L_{O}(T)-L_{W}(t)$ would also require a significant prior likelihood $L_{O}(t)$ for any reasonably small values of the parameter $\alpha$.
@@ 3 Strong Priors in Generative Models
The term “strong priors" in this context was introduced by McKenzie et al. (2023) as one of the categories of LM tasks that exhibit inverse scaling trends. The difficulty of this category of tasks seems to largely stem from local context overly influencing the predicted continuation. We aim to extend the informal characterisation in McKenzie et al. (2023) with a more comprehensive definition.
@@
One example of the phenomenon is an autoregressive language model has been prompted to write a common sequence (the alphabet) with a variation in the middle (a skipped letter). We empirically observe that once the LM has written enough of the common sequence, it fails to add the variation. We hypothesise that the model is overgeneralising from a frequent pattern in its training data, ignoring other context.
@@
In instances where $p_{M}(t|TD)\approx p_{M}(t|D)$, the difference between the two distributions may hold significant information: the correct continuation is often more probable in $p_{M}(t|TD)$ than $p_{M}(t|D)$. We present a novel method to compensate for strong local priors based on our model in Section 5.
There also exist non-local strong priors, i.e. cases where $T$ and $D$ are intertwined in more complex way but still allow for a meaningful systematic distinction between task and data. We mostly focus on the local case, but the above foundational definitions as well as the proposed method can be adapted for such non-local scenarios. We include a concrete example in Appendix A and empirical measurements of $\delta$ in Appendix C.
@@ 4 Language Models as Products of Experts
In this section, we delve into our proposed model that addresses the challenge of strong priors in generative language models. Our approach models the generative distribution as a product of experts (i.e. a geometric mixture) of two components: one that captures the correct continuation of context, and another that captures continuation based on strong local priors. We then define a new distribution that weighs between these components to produce more contextually appropriate completions.
@@
Geometric Mixture Model. We assume that $p_{M}$ can be modeled as a geometric mixture of $p_{C}$ and $p_{L}$, also known as Product of Experts:
@@
Approximating a strong prior model. To construct a strong prior model $L$ an approximation of model $M$ that is more prone to be influenced by strong priors - we alter the prompts to accentuate this behavior, and then employ $M$ on these modified prompts. The main modification we consider is stripping the task from the prompt, although in Appendix B we discuss a different approach based on comparing model outputs with and without a weakening system prompt that is task-agnostic.
@@
For instance, for the task: "View number as text. Do not perform computation. Q: What is the first digit of 23+63?", $P$ represents the text: "View number as text. Do not perform computation." and $t_{1}\dots t_{n}$ captures " Q: What is the first digit of 23+63?".

### v3

@@ 1 Introduction
Language models (LMs) have come to occupy a central role in a wide range of tasks, from data processing to the creation of instruction-following assistants. These models seem to both increase in performance and also develop new capabilities as they scale up in parameters, especially when given examples of a task (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). They see both widespread public use, and increasing integration into sensitive tasks where their versatile capabilities are key (Javaid et al., 2022). However, this increasing reliance on LMs raises concerns about their reliability and potential vulnerabilities. Their extremely general ability to recognise patterns may come with a cost: language models are susceptible to distractor tasks, unintended secondary tasks which a model can infer from the prompt in addition to the intended one, leading to unwanted outputs (Wei et al., 2023).
Distractor tasks include both semantically complex cases like prompt injection, in which the distractor task of following the most recently injected instruction overrides the intended task of following the initial instructions, and much more semantically simple cases where a model defaults to outputting a common pattern regardless of prior context (Greshake et al., 2023; McKenzie et al., 2023).
@@
Even cutting-edge models struggle with trivial tasks in the presence of very common patterns. For example, as of September 2024 GPT-4o, when prompted with ”Write all the numbers up to 30 except multiples of 13, and nothing else”, will correctly exclude 13 but fail to exclude 26. Other examples of such distractor tasks include repeating common misconceptions instead of true answers (Lin et al., 2022) and failing to write mathematical proofs when the correct proof is very similar to an incorrect but much more common proof (Collins et al., 2024).
@@
Furthermore, for some tasks of this type, larger models actually perform worse: similar models with more parameters can be more likely to repeat common misconceptions and succumb to prompt injection attacks (McKenzie et al., 2023; Lin et al., 2022). This phenomenon, known as ‘inverse scaling’ (Lin et al., 2022), reveals a pressing need to better understand how these problems arise and how they can be fixed. Unlike other problems with LMs which are usually resolved by scaling up the model and training data size, the existing problems might become even more pronounced, and other similar issues might emerge.
1.1 Prior-Aware Decoding
To better understand and address the problem of distractor tasks, we characterise language models as a product of experts (Hinton, 1999) which combine the predictions of several component models, each component model predicting the next token assuming a particular task or context. This framework, presented in detail in Section 4, allows us to decompose the model’s behavior into distinct ”experts” that may be responsible for different aspects of the output, which in turn lets us interpret the problem of distractor tasks as models overweighing a particular expert.
Using this framework, we present a technique to reduce the influence of distractor tasks in cases where the prompt can be divided into one part containing the intended task and one part containing the distractor task and other important information. In this case, we can contrast the outputs the model gives for the full prompt with the outputs it gives for just the part which yields the distractor task, extrapolating to infer how it would behave with less influence from the distractor task. This is a form of contrastive decoding (Li et al., 2023). Our approach differs from previous methods by specifically targeting the influence of distractor tasks and leveraging the product of experts model
More specifically, our method generates the original logits $L_{O}$ from the original prompt, as well as the weakened logits $L_{W}$ from the portion of the prompt which does not contain the original task. Even though $L_{O}$ and $L_{W}$ would both, by default, favour undesired outputs consistent with the distractor task, $L_{O}$ gives a higher probability to outputs consistent with the intended task. So, by sampling instead from a linear combination of the logits, $L_{O}+\alpha(L_{O}-L_{W})$ we can selectively favour outputs which are more consistent with the original task. ($\alpha$ here is a tunable hyperparameter.) Using this technique we can mitigate the influence of the distractor task, as we illustrate in Figure 1.
@@ 1.1 Prior-Aware Decoding
Figure 1: Illustration of our geometric mixture model, also referred to as product of experts, and Prior-Aware Decoding in a task requiring a modification of a very common sequence.
@@
Our primary baseline is the unmodified completion according to the same model. Across the 44 model-task combinations we find that our technique outperforms the baseline on 41 combinations, with a median absolute increase in proportion of tasks completed of 40%. See Figure 2 and Section 6 for the results. These improvements demonstrate the potential of our method to significantly enhance the reliability and performance of language models across a variety of model architectures in scenarios featuring distractor tasks.
The technique requires two choices: selection of the weakened prompt, and the coefficient $\alpha$. We offer consistent approaches to generating the weakened prompts for each of our datasets and also propose a general method for weakened prompt construction. We explore a range of values for $\alpha$. Crucially, this method does not require retraining: it operates at inference time by making two queries instead of one and computing a linear combination of the resulting logits.
The improvements in accuracy due to our technique suggest that viewing LMs as product of experts models can be a useful framework for understanding and improving the behaviour of these models. Our research provides new approaches for utilizing capabilities which are dormant in models and which have been hard to elicit with prompting alone. This includes capabilities of immediate practical importance, such as resistance to prompt injection attacks, as well as understanding how to correct for biases learned in the training data.
@@
  • Interpreting language models as geometric mixtures (i.e. a product of experts) to develop a model of how distractor tasks cause poor performance on certain inputs.
@@ 2 Related Work
Inverse Scaling. For most tasks, performance scales together with the model scale (model size, dataset size, and training compute). However, certain tasks exhibit inverse scaling, where increasing the scale of the model leads to worse performance. The term was introduced by Lin et al. (2022) to characterise the tendency of LMs to repeat common misconceptions. McKenzie et al. (2023) further developed the notion and provided the Inverse Scaling dataset as well as several informal categories. Wei et al. (2023) adds to the discussion, showing that of the original eleven tasks in the Inverse Scaling dataset, only four persist in demonstrating inverse scaling on models of up to 540B parameters. We apply our method to three of those four tasks. Our work is the first to offer a formal account of the mechanisms that have so far only been informally hypothesised in Wei et al. (2023), and we use this to develop a novel technique that improves performance.
@@
Problems Related to Strong Priors. While in this paper we investigate the problem of strong priors in the domain of language modelling (discussed primarily in Section 3), a somewhat similar problem also occurs in diffusion models for image generation. Diffusion models have been observed to sometimes struggle to generate images that have a precise correspondence to the prompt. This problem is especially pronounced for complex prompts that include rare words, or describe implausible scenes, spatial relations, or compositionality. One technique for mitigating this problem in diffusion models is classifier-free guidance (Ho & Salimans, 2021). It samples the diffusion model using a linear combination of conditioned and unconditioned score estimates $(1+w)\cdot e(z_{t},c)-w\cdot e(z_{t})$, where $z_{t}$ is the image at the denoising step $t$, $c$ is the conditioning information such as the prompt, $e$ is the model and $w$ is the guidance weight. High guidance weight improves text-image alignment (Saharia et al., 2022), although high values often lead to over-saturated images or have a generally unnatural appearance. The technique that we present in this paper uses a somewhat similar combination of differently conditioned model outputs to mitigate the problem of strong priors in LMs.
@@
Surface Form Competition and DC-PMI. Holtzman et al. (2021) coin ‘surface form competition’ as a phenomenon where different tokens representing the same concept compete for probability. For instance, the terms “computer” and “PC” represent the same concept but are different surface forms. Since the probability mass is finite, this competition lowers the probability of the correct answer. Domain Conditional Pointwise Mutual Information (DC-PMI) addresses this issue by reweighing each option according to its likelihood within the context of a specific task, calculated by sampling from a subset of the prompt. This approach yields gains in zero-shot performance for a variety of GPT-2 and GPT-3 models on a selection of multiple-choice datasets. Our work applies a similar kind of technique - reweighing based on likelihood - but we do so to address a different set of tasks, including those which exhibit inverse scaling.
@@
Contrastive Inference Methods. Contrastive Inference methods steer language models towards desired outputs by taking multiple samples across different settings and favouring outputs which are comparatively more likely in one sample than in another. The original implementation in Li et al. (2023) takes output distributions using the same prompt in two models, one smaller and one larger. It then contrasts the logits from the two models, selecting the token with the largest difference between the original ($L_{O}$) and smaller model ($L_{S}$), i.e. $\max_{t\in T^{\prime}}(L_{O}(t)-L_{S}(t))$, limited to a subset $T^{\prime}$ of top tokens to mitigate noise in low-likelihood tokens.
Our approach is most similar to the subsequent techniques introduced in Shi et al. (2024) and Malkin et al. (2022), which apply contrastive decoding to pairs of shorter and longer prompts, using the same formula we apply. Malkin et al. (2022) contrasts the full prompt against only the text generated so far to increase long-range semantic coherence, finding improvement in generic language understanding and knowledge retrieval tasks. Shi et al. (2024) uses an equivalent formula to handle conflicts between prior knowledge and information in the context. They also interpret the combination as a product-of-experts.
Compared to these works, ours frames the language model itself as a product-of-experts that must select between specific competing tasks, some of which are unwanted distractors, and the resulting manipulation of logits as a reweighing of the experts. Furthermore, while Malkin et al. (2022) and Shi et al. (2024) focus on models failing to account for knowledge and facts, we emphasise how they can include even extremely semantically simple patterns like sequences of numbers. We also draw the connection to inverse scaling.
@@ 3 Strong Priors in Generative Models
The term “strong priors” in this context was introduced by McKenzie et al. (2023) as one of the categories of LM tasks that exhibit inverse scaling trends. The difficulty of this category of tasks seems to largely stem from local context overly influencing the predicted continuation. We aim to extend the informal characterisation in McKenzie et al. (2023) with a more comprehensive definition.
@@
One example of the phenomenon is an autoregressive language model that has been prompted to write a common sequence (the alphabet) with a variation in the middle (a skipped letter). We empirically observe that once the LM has written enough of the common sequence, it fails to add the variation. We hypothesise that the model is overgeneralising from a frequent pattern in its training data, ignoring other context.
@@
In instances where $p_{M}(\cdot|TD)\approx p_{M}(\cdot|D)$, the difference between the two distributions may hold significant information: the correct continuation is often more probable in $p_{M}(\cdot|TD)$ than $p_{M}(\cdot|D)$. We present a novel method to compensate for strong local priors based on our model in Section 5.
While we focus on the local case, the above foundational definitions as well as the proposed method can be adapted for non-local scenarios where $T$ and $D$ are intertwined in more complex way but still allow for a meaningful systematic distinction between task and data. We include a concrete example in Appendix A and empirical measurements of $\delta$ in Appendix D.
@@ 4 Language Models as Products of Experts
In this section, we present our proposed model for addressing the challenge of distractor tasks emerging from strong priors in generative language models. Our approach models the generative distribution as a product of experts (i.e. a geometric mixture) of two components: one that captures the correct continuation of context based on the intended task, and another that captures continuation based on the distractor task induced by strong local priors. We then define a new distribution that weighs between these components to produce more contextually appropriate completions.
@@
Geometric Mixture Model. We assume that $p_{M}$ can be modeled as a geometric mixture of $p_{C}$ and $p_{L}$, also known as a Product of Experts:
@@
Approximating a strong prior model. To construct a strong prior model $L$ an approximation of model $M$ that is more prone to be influenced by strong priors - we alter the prompts to accentuate this behavior, and then employ $M$ on these modified prompts. The main modification we consider is stripping the task from the prompt, although in Appendix C we discuss a different approach based on comparing model outputs with and without a weakening system prompt that is task-agnostic.
@@
For instance, for the task: ”View number as text. Do not perform computation. Q: What is the first digit of 23+63?”, $P$ represents the text: ”View number as text. Do not perform computation.” and $t_{1}\dots t_{n}$ captures ” Q: What is the first digit of 23+63?”.